# Automated Structural-level Alignment of Multi-view TLS and ALS Point Clouds in Forestry

### v3

@@ Abstract
Access to highly detailed models of heterogeneous forests, spanning from the near surface to above the tree canopy at varying scales, is increasingly in demand. This enables advanced computational tools for analysis, planning, and ecosystem management. LiDAR sensors, available through terrestrial (TLS) and aerial (ALS) scanning platforms, have become established as primary technologies for forest monitoring due to their capability to rapidly collect precise 3D structural information directly. Selection of these platforms typically depends on the scales (tree-level, plot, regional) required for observational or intervention studies. Forestry now recognizes the benefits of a multi-scale approach, leveraging the strengths of each platform while minimizing individual source uncertainties. However, effective integration of these LiDAR sources relies heavily on efficient multi-scale, multi-view co-registration or point-cloud alignment methods. In GPS-denied areas, forestry has traditionally relied on target-based co-registration methods (e.g., reflective or marked trees), which are impractical at scale. Here, we propose an effective, target-less, and fully automatic co-registration method for aligning forest point clouds collected from multi-view, multi-scale LiDAR sources. Our co-registration approach employs an incremental alignment strategy, grouping and aggregating 3D points based on increasing levels of structural complexity. This strategy aligns 3D points from less complex (e.g., ground surface) to more complex structures (e.g., tree trunks/branches, foliage) sequentially, refining alignment iteratively. Empirical evidence demonstrates the method’s effectiveness in aligning TLS-to-TLS and TLS-to-ALS scans across various ecosystem conditions, including pre/post fire treatment effects. In TLS-to-TLS scenarios, rotation and translation parameter RMSE errors were $<0.1^{\circ}$ and $<2$ cm, respectively. For TLS-to-ALS, corresponding errors were $<0.7^{\circ}$ and $\leq 6$ cm. To the best of our knowledge, our method is the first capable of automatically co-registering forest point clouds from both TLS-to-TLS and TLS-to-ALS scenarios without relying on targets, while achieving high performance.
@@ 1 Introduction
Remote sensing techniques, including terrestrial laser scanning (TLS), aerial laser scanning (ALS), and optical photogrammetry, have been established as technologies offering time and cost economies at scale in forest surveys [3]. Among these, LiDAR-based scanning has gained widespread interest due to its inherent capability to rapidly collect precise 3D structural information over large regional extents [4]. Airborne based LiDAR for example, utilizing accurate position sensors including RTK, enables data collection at the regional/landscape scale. It has been shown to be effective for forest characterization and prediction of forest factor metrics [5]. However, some limitations exist as it has difficulty penetrating the tree canopy cover, resulting in low-resolution spatial variability along the tree stand direction, even with its powerful but computationally demanding full-waveform (FW) based technology. In-situ TLS, on the other hand, is complementary to ALS as it is capable of collecting information at a much higher 3D spatial resolution along the full vertical tree stand direction (i.e., ground surface, mid-tree story, and canopy vegetation) [6]. However, TLS is limited to a few tens of meters in range, and scaling its advantages over larger area extents requires more sophisticated methods of sensor placement stratification and data source fusion. Having access to a diversity of sensing sources and a proper way to combine them can enhance capabilities, enabling precise 3D information from near-surface vegetation to the top of the canopy at scale—a characteristic rarely available when a single scanning mode is used [6]. Here, we focus on this problem and propose a new 3D point cloud co-registration method suitable for forest environments, with the purpose of combining multi-view, multi-source LiDAR data. The proposed method is fully automatic, does not require any targets, and does not rely on the presence of man-made objects in the scene, as is typically required by conventional point cloud co-registration approaches in forestry.
@@
In light of this discussion, we propose a new method to co-register point clouds from forest ecosystems that is fully automatic and does not require any manual target placements. The proposed method finds the co-registration parameters by using a notion of 3D structural complexity that encodes information compactness or compression about the point cloud composition. Here, we use a dense 3D plane approximation to the point cloud and a mixture of von-Mises Fisher (vMF) distributions to represent and extract the point cloud composition and the complexity or informativeness of its components. We show that grouping and matching LiDAR scan point subsets based on this notion of 3D structural complexity is an effective strategy for co-registering pairs of scans from highly complex and heterogeneous forest environments and under a variety of conditions. The intuition behind our incremental co-registration strategy relies on the fact that forests are composed of 3D structures of varying complexity (e.g., ground, tree trunks/branches, foliage), and designing alignment approaches should exploit those that are less complex (e.g., ground, tree trunks) to find a rough alignment first, and then use those of increasing complexity (e.g., foliage) as a refinement. Figure 1 illustrates a preliminary example aligning five plot-scale, multi-view TLS scans and a regional ALS scan collected from a New Mexico, USA forest. Advantages of the proposed method include its effectiveness in registering LiDAR scans in complex forest environments without requiring any position sensors or reference targets, explicit extraction of trees or other semantic features, while also being fully automatic and scalable to large area extents with the ability to co-register many TLS-to-TLS and TLS-to-ALS scans. In the following, Section 2 describes our approach, while Section 3 empirically demonstrates the overall performance of our method. Finally, Section 4 includes a short discussion about the benefits of the proposed approach in the context of forestry, and Section 5 summarizes our contributions.
@@ 2.1 Pair-wise co-registration
where the first summation is over the number of $K$ complexity groups, $H(\tilde{\mathbf{X}}^{\{\mathbf{Y}==k\}}|~{}\boldsymbol{\mu}_{k},\boldsymbol{\kappa}_{k})$ is the conditional entropy of the $k^{\text{th}}$ vMF component given parameters $\boldsymbol{\mu}_{k},\boldsymbol{\kappa}_{k}$, and the second summation on the right hand side is over the 3D planes associated with each group $k$. In other words, the overall structural complexity of a point cloud $\mathbf{X}$ can be described by the summation over the individual component complexities $\text{SC}(\tilde{\mathbf{X}},k)$. The intuition behind Eq.(2) is that 3D points corresponding to different components of the forest exhibit varying levels of information compactness or concentration. For instance, points on the forest ground surface are expected to have high concentration or information compactness in their 3D plane approximations, indicating low structural complexity. On the other hand, points representing foliage are expected to have low concentration and more variability in their 3D plane orientations, suggesting high structural complexity. It is worth noting that notions of structural scene complexity have been utilized previously (e.g., in [38, 39]) to sub-sample scenes at different complexity scales based on waveform envelope shapes in full-waveform LiDAR data.
@@ 2.1.3 Incremental Alignment Optimization
Algorithm 1 Pair-wise multi-structural level co-registration
@@ 3 Experiments
In the following set of experiments, validation is conducted on real LiDAR point cloud datasets collected from various forest ecosystems. Specifically, the experiments include: (1) multi-view TLS scans, and (2) co-located TLS and ALS scans. The datasets were primarily sourced from the Bandelier National Monument in New Mexico, USA, encompassing high-elevation mixed-conifer forests, and from Ft. Stewart, GA, USA, unless otherwise noted. The implementations were carried out in Python, utilizing the VTK-based PyVista module for visualization, the scikit-learn implementation of the mixture of vMF algorithms [36], and selected tools from Open3D [42], including ICP and downsampling functionalities.
@@ 3.1 Datasets
Performance Metrics: All quantitative evaluations in our study were conducted using the registration root mean square error (RMSE) metric. The registration RMSE was measured relative to the ground truth six degrees of freedom orientation and position registration parameters. Ground truth data varied depending on the experiment. In Subsections 3.2.1 to 3.2.2, ground truth was established by running our approach on TLS ground scans for 10 trials per pairwise scan, averaging registration parameters, and ensuring consistency with augmented ALS. Consistency with ALS was confirmed by comparing relative orientations and positions derived from TLS scan centers projected onto the ground plane in ALS scans. It’s noteworthy that ALS aligned with RTK provides accuracies at the centimeter level. Following these steps, the resulting registrations were validated by a trained expert. In Section 3.2.3, ground truth was obtained from a reference-based method using a LiDAR dataset collected in a forest ecosystem. Evaluations were also benchmarked against other well-known standard co-registration methods. Overlap was quantified as the percentage of points in the source scan having corresponding points in the target scan within a distance of 25 cm, relative to the total number of points in the pair of scans. For registration involving multiple scans, overlap was similarly calculated for every pairwise combination within a plot. Finally, inlier RMSE measures the minimum RMSE distance between 3D points in aligned source and target point clouds that are at least 25 cm apart. This metric provides a quantitative assessment of the proximity between aligned source and target point clouds.
@@ 3.2.1 Overall results
The pairwise registration method described in Section 2.1 was evaluated on pairs of multi-view TLS point clouds. Example results are illustrated in Fig. 3. The first row shows four views of the registration of two color-coded scans positioned 15 x 5 meters apart in a forest site featuring white fir, limber pine, aspen, Douglas fir, and surface bark, trunk, and bushes. Fig. 3(a) depicts an initial example where forest features are misaligned. In contrast, Figs. 3(b-d) display different views of the alignment results after applying our co-registration approach. The second row presents qualitative results demonstrating the alignment among five scans within a single plot using our pairwise co-registration method relative to the center point-cloud scan. Overall, Fig. 3 shows well-aligned point clouds with consistent object alignment across scan transitions. Notably, we utilized three structural complexity levels ($K=3$) in the mixture of vMF distributions and empirically observed that the resulting groups mainly corresponded to: (1) the ground surface, (2) tree trunks/branches, and (3) foliage. This finding remained consistent across various cases of forest structure captured with TLS.
@@
Figure 4: Robustness of the structural complexity-based algorithm against random alignment initializations and forest heterogeneity in the TLS-to-TLS co-registration problem.
To comprehensively evaluate the performance of our proposed approach across diverse forest conditions encompassing heterogeneous vegetation and varying topographical gradients—we conducted pairwise registrations in 21 distinct inventory plots, totaling 84 pairs of scans. Initially, each scan, was synthetically misaligned by randomly and uniformly perturbing all six co-registration parameters simultaneously. These perturbations were constrained within ±$\pm 45^{\circ}$ degrees for rotations and ±$\pm 15$ meters for translations relative to the ground truth alignment. We performed 10 trials for each pair of scans (resulting in 840 co-registrations in total), each trial using a different initialization drawn from a uniform distribution. The performance of our co-registration approach across 840 co-registrations is depicted in Figure 4. Each point on the horizontal axis represents one trial in one of the 84 pairs of scans evaluated. The vertical axis shows the error, measured in meters or degrees, across all 10 trials for each alignment parameter (rotation and translation). The initializations are represented in blue, indicating large but random initial errors, while the final errors after co-registration are shown in orange, demonstrating successful reduction in alignment error for all six parameters. The overall RMSE values for each co-registration parameter, computed across all scan pairs and initializations, are summarized in the second row of the legend within each plot. We found that the RMSE is less than $0.1^{\circ}$ for rotation and less than $0.02$(m.) meters for translation parameters. These results confirm that the method is flexible and robust, capable of handling vegetation diversity, topographical variations, and random alignment initializations effectively.
@@ 3.2.2 Non-uniform ground terrain
Fig. 6 provides a summary of our experiments evaluating registration performance as a function of percent overlap. Each plot corresponds to an alignment parameter, with blue and orange shapes representing the distributions of initialization and estimate errors relative to ground truth, respectively. The width of each distribution indicates the frequency distribution of error values (vertical axis), while the horizontal axis shows the percent overlap tested across four settings: $28,15,5,1\%$. A total of 1000 co-registration trials per percent overlap were conducted, resulting in 4000 co-registrations in total. For each trial, we randomly initialized the six alignment parameters (simultaneously randomized) from a uniform distribution centered at the true alignment, deviating by ±$\pm 45^{\circ}$ in rotation and ±$\pm 15$ m in translation (as shown on the vertical axis). In Fig.6, it is evident that errors for all six parameters remain concentrated around zero, regardless of initialization and the percent overlap tested. Table 1 summarizes the quantitative RMSE values, providing a quantification of this error variability across different alignment parameters and percent overlaps. The consistently small RMSE values across varying levels of overlap underscore the effectiveness and robustness of our approach, even under conditions of limited point-cloud overlap.
@@
We compared the performance of our pair-wise registration approach against a target-based method using Styrofoam spheres as benchmarks. TLS scans from the dataset in [43] were registered using our approach and compared with the target-based method, which provided registration parameters already validated as ground truth in the referenced dataset. This dataset consists of scans from 17 plots, each with 4 scans per plot arranged in a triangular pattern (one at the center and three at each corner 15 meters apart from the center), along with 9 manually placed targets within the scene.
We conducted a total of 100 trials per available scan pair, initializing each trial uniformly with random rotations and translations within ±$\pm 45^{\circ}$ in rotation and ±$\pm 15$ m in translation error relative to ground truth. Table 2 summarizes the RMSE for all 51 pairs of scans across 100 trials each, relative to the registration parameters obtained from the target-based method, computed for each plot relative to its center scan.
Table 2: Registration RMSE relative to benchmark target-based method
@@
0.058 0.065 0.032 0.016 0.012 0.007
Based in Table 2, our framework achieves accuracies comparable to those of target-based methods. This validates that our approach performs equivalently to target-based methods, while offering the advantages of full automation without the need for manually placed targets or semantic object detection.
@@
Additional experiments were conducted to compare the performance of our co-registration method against standard methods. Specifically, we selected automatic, targetless methods that do not rely on lines, contours, planes, normals, or edge features, which are typically assumed by methods designed for urban scenarios. The methods included iterative closest point [44], coherent point drift [45], gaussian mixture models [46], support vector registration [47], trees of gaussian mixtures [48], filterReg [49], and Bayesian coherent point drift [50], all available in the Probreg Python package [51]. For this comparison, we randomly selected 50 pairs of point clouds from the available forest plots in the NM dataset. Each pair was tested using 100 trials with random initialization from the true co-registration parameters, deviating by ±$\pm 30$ degrees and ±$\pm 15$ meters in a uniform distribution. The RMSE relative to the six degrees of freedom was then computed and summarized in Table 3.
Table 3: Methods comparison of co-Registration performance by RMSE

### v4

# Juan Castorena L. Turin Dickman Adam J. Killebrew James R Gattiker Rod Linn E. Louise Loudermilk Los Alamos National Laboratory, Los Alamos, NM, 48124 USA Southern Research Station, 200 W.T. Weaver Blvd. Asheville, NC 28804-3454, USA

@@ Abstract
Access to highly detailed models of heterogeneous forests, spanning from the near surface to above the tree canopy at varying scales, is increasingly in demand. This enables advanced computational tools for analysis, planning, and ecosystem management. LiDAR sensors, available through terrestrial (TLS) and aerial (ALS) scanning platforms, have become established as primary technologies for forest monitoring due to their capability to rapidly collect precise 3D structural information directly. Selection of these platforms typically depends on the scales (tree-level, plot, regional) required for observational or intervention studies. Forestry now recognizes the benefits of a multi-scale approach, leveraging the strengths of each platform while minimizing individual source uncertainties. However, effective integration of these LiDAR sources relies heavily on efficient multi-scale, multi-view co-registration or point-cloud alignment methods. In GPS-denied areas, forestry has traditionally relied on target-based co-registration methods (e.g., reflective or marked trees), which are impractical at scale. Here, we propose ForestAlign: an effective, target-less, and fully automatic co-registration method for aligning forest point clouds collected from multi-view, multi-scale LiDAR sources. Our co-registration approach employs an incremental alignment strategy, grouping and aggregating 3D points based on increasing levels of structural complexity. This strategy aligns 3D points from less complex (e.g., ground surface) to more complex structures (e.g., tree trunks/branches, foliage) sequentially, refining alignment iteratively. Empirical evidence demonstrates the method’s effectiveness in aligning TLS-to-TLS and TLS-to-ALS scans locally, across various ecosystem conditions, including pre/post fire treatment effects. In TLS-to-TLS scenarios, parameter RMSE errors were less than 0.75 degrees in rotation and 5.5 cm in translation. For TLS-to-ALS, corresponding errors were less than 0.8 degrees and 8 cm, respectively. These results, show that our ForestAlign method is effective for co-registering both TLS-to-TLS and TLS-to-ALS in such forest environments, without relying on targets, while achieving high performance.
@@ 1 Introduction
Remote sensing techniques, including terrestrial laser scanning (TLS), aerial laser scanning (ALS), and optical photogrammetry, have been established as technologies offering time and cost economies at scale in forest surveys [3]. Among these, LiDAR-based scanning has gained widespread interest due to its inherent capability to rapidly collect precise 3D structural information over large regional extents [4]. Airborne based LiDAR for example, utilizing accurate position sensors including RTK, enables data collection at the regional/landscape scale. It has been shown to be effective for forest characterization and prediction of forest factor metrics [5]. However, some limitations exist as it has difficulty penetrating the tree canopy cover [6], resulting in low-resolution spatial variability along the tree stand direction, even with its powerful but computationally demanding full-waveform (FW) based technology. In-situ TLS, on the other hand, is complementary to ALS as it is capable of collecting information at a much higher 3D spatial resolution along the full vertical tree stand direction (i.e., ground surface, mid-tree story, and canopy vegetation) [7]. However, TLS is limited to a few tens of meters in range, and scaling its advantages over larger area extents requires more sophisticated methods of sensor placement stratification and data source fusion. Having access to a diversity of sensing sources and a proper way to combine them can enhance capabilities, enabling precise 3D information from near-surface vegetation to the top of the canopy at a characteristic scale rarely available when a single scanning source is used [7]. Here, we focus on this problem and propose a new 3D point cloud co-registration method suitable for forest environments, with the purpose of combining multi-view, multi-source LiDAR data. The proposed method is fully automatic, does not require any targets, and does not rely on the presence of man-made objects in the scene, as is typically required by conventional point cloud co-registration approaches in forestry.
@@
In light of this discussion, we propose a new method termed automatic forest structure-based alignment (ForestAlign) to co-register point clouds from forest ecosystems that is fully automatic and does not require any manual target placements. The proposed method finds the co-registration parameters by using a notion of 3D structural complexity that encodes information compactness or compression about the point cloud composition. Here, we use a dense 3D plane approximation to the point cloud and a mixture of von-Mises Fisher (vMF) distributions to represent and extract the point cloud composition and the complexity or informativeness of its components. We show that grouping and matching LiDAR scan point subsets based on this notion of 3D structural complexity is an effective strategy for co-registering pairs of scans from highly complex and heterogeneous forest environments and under a variety of conditions. The intuition behind our incremental co-registration strategy relies on the assumption that forests are composed of underlying 3D structures of varying complexity (e.g., ground, tree trunks/branches, foliage), and designing alignment approaches should exploit those that are less complex (e.g., ground, tree trunks) to find a rough alignment first, and then use those of increasing complexity (e.g., foliage) as a refinement. Figure 1 illustrates a preliminary example aligning five plot-scale, multi-view TLS scans and a regional ALS scan collected from a New Mexico, USA forest. Advantages of the proposed method include its effectiveness in registering LiDAR scans in complex forest environments without requiring any position sensors or reference targets, explicit extraction of trees or other semantic features, while also being fully automatic and scalable to large area extents with the ability to co-register many TLS-to-TLS and TLS-to-ALS scans. In the following, Section 2 describes our approach, while Section 3 empirically demonstrates the overall performance of our method. Finally, Section 4 includes a short discussion about the benefits of the proposed approach in the context of forestry, and Section 5 summarizes our contributions.
@@ 2.1 Pair-wise co-registration
where the first summation is over the number of $K$ complexity groups, $H(\tilde{\mathbf{X}}^{\{\mathbf{Y}==k\}}|~{}\boldsymbol{\mu}_{k},\boldsymbol{\kappa}_{k})$ is the conditional entropy of the $k^{\text{th}}$ vMF component given parameters $\boldsymbol{\mu}_{k},\boldsymbol{\kappa}_{k}$, and the second summation on the right hand side is over the 3D planes associated with each group $k$. In other words, the overall structural complexity of a point cloud $\mathbf{X}$ can be described by the summation over the individual component complexities $\text{SC}(\tilde{\mathbf{X}},k)$. For instance, points on the forest ground surface are expected to have high concentration or information compactness in their 3D plane approximations, indicating low structural complexity. On the other hand, points representing foliage are expected to have low concentration and more variability in their 3D plane orientations, suggesting high structural complexity. It is worth noting that notions of structural scene complexity have been utilized previously (e.g., in [39, 40]) to sub-sample scenes at different complexity scales based on waveform envelope shapes in full-waveform LiDAR data.
@@ 2.1.3 Incremental Alignment Optimization
Algorithm 1 Automatic Forest structure-based alignment (ForestAlign)
@@ 3 Experiments
In the following set of experiments, validation is conducted on real LiDAR point cloud datasets collected from various forest ecosystems. Specifically, the experiments include co-registration of: (1) multi-view TLS scans, and (2) co-located TLS and ALS scans. The datasets were primarily sourced from the Bandelier National Monument in New Mexico, USA, encompassing high-elevation mixed-conifer forests, and from Ft. Stewart, GA, USA, unless otherwise noted. The implementations were carried out in Python, utilizing the VTK-based PyVista module for visualization, the scikit-learn implementation of the mixture of vMF algorithms [37], and selected tools from Open3D [43], including ICP and downsampling functionalities.
@@ 3.1 Datasets
Performance Metrics: All quantitative evaluations in our study were conducted using the registration root mean square error (RMSE) metric. The registration RMSE was measured relative to the ground truth six degrees of freedom orientation and position registration parameters. Ground truth was obtained, whenever additional source data was available, by cross-validation.
In cases of overlapping TLS-to-TLS pair-wise co-registration, the ground truth parameters were obtained by first manually co-registering each TLS scan in the pair to the overlapping ALS data (note that, for computational efficiency, only a truncated ALS point cloud was used, not the entire regional ALS). After deriving the relative TLS-to-ALS co-registration for each scan in the pair, we calculated the relative position between the TLS scans in the ALS coordinate system by applying the TLS-to-ALS transformations. We then transformed this relative position back into the target TLS reference coordinate system using the inverse of the TLS-to-ALS transformation for the target scan.
In the TLS (source)-to-ALS (target) co-registration case, a similar process was followed to obtain ground truth. Here, pairs of overlapping TLS-to-TLS and corresponding truncated ALS point clouds are used. The truncated ALS removes 3D points outside a 50 × 50 meter area in the lat/lon direction, centered at a TLS GPS reference location. First, manual co-registration between TLS (source)-to-TLS (cross) was performed, and then manual co-registration between TLS (cross)-to-ALS was obtained. This process provides a transformation from the source TLS scan being co-registered to the ALS scan, on which we also apply ICP using the same settings as in the TLS-to-TLS case. When no additional overlap between TLS-to-TLS or TLS-to-ALS is available (as in the Ft. Stewart dataset and reference target datasets), cross-validation is not used. Instead, we proceeded directly to manual co-registration between the involved TLS (source)-to-ALS (target) or TLS (source)-to-TLS (target) pairs.
The manual co-registration method we used is from the open-source CloudCompare software [44], which provides interactive visualizations to manually select corresponding points between point clouds. Given the corresponding points, the software automatically co-registers the pairs of point clouds. Our methodology involved selecting 5 corresponding points in the source and target scans, distributed across the entire point cloud. This distribution constrains the alignment more effectively than selecting points that are closely clustered together. It is important to note that 5 corresponding points exceed the minimum requirement of 3 for the Procrustes algorithm [45] to perform co-registration. Additionally, after manual co-registration, we apply ICP with a downsampling factor of 0.015 m and a closest point threshold of 0.25 m over 50 iterations to further refine the ground truth estimates from the manual alignment.
In Section 3.2.3, ground truth was obtained from a reference-based method using a LiDAR dataset collected in a forest ecosystem. Evaluations were also benchmarked against other well-known standard co-registration methods. Overlap was quantified as the percentage of points in the source scan having corresponding points in the target scan within a distance of 25 cm, relative to the total number of points in the pair of scans. For registration involving multiple scans, overlap was similarly calculated for every pairwise combination within a plot. Finally, inlier RMSE measures the minimum RMSE distance between 3D points in aligned source and target point clouds that are at least 25 cm apart. This metric provides a quantitative assessment of the proximity between aligned source and target point clouds.
@@ 3.2.1 Overall results
The pairwise registration method described in Section 2.1 was evaluated on pairs of multi-view TLS point clouds. Example results are illustrated in Fig. 3. The first row shows four views of the registration of two color-coded scans positioned 15 x 5 meters apart in a forest site featuring white fir, limber pine, aspen, Douglas fir, and surface bark, trunk, and bushes. Fig. 3(a) depicts an initial example where forest features are misaligned. In contrast, Figs. 3(b-d) display different views of the alignment results after applying our co-registration approach. The second row presents qualitative results demonstrating the alignment among five scans within a single plot using our pairwise co-registration method relative to the center point-cloud scan. Overall, Fig. 3 shows well-aligned point clouds with consistent object alignment across scan transitions. Notably, we utilized three structural complexity levels ($K=3$) in the mixture of vMF distributions and empirically observed that the resulting groups mainly corresponded to: (1) the ground surface, (2) tree trunks/branches, and (3) foliage. This finding remained qualitatively consistent across various cases of forest structure captured with TLS.
@@
Figure 4: Robustness of the structural complexity-based algorithm against random alignment initializations in the TLS-to-TLS co-registration problem
To comprehensively evaluate the performance of our proposed approach across diverse forest conditions encompassing heterogeneous vegetation and varying topographical gradients—we conducted pairwise registrations in 21 distinct inventory plots, totaling 84 pairs of scans. Initially, each scan, was synthetically misaligned by randomly and uniformly perturbing all six co-registration parameters simultaneously. These perturbations were constrained within less than ±$\pm 45^{\circ}$ degrees for rotations and ±$\pm 15$ meters for translations relative to the ground truth alignment. We performed 10 trials for each pair of scans (resulting in 840 co-registrations in total), each trial using a different initialization drawn from a uniform distribution. The performance of our co-registration approach across 840 co-registrations is depicted in Figure 4. Each point on the horizontal axis represents one trial in one of the 84 pairs of scans evaluated. The vertical axis shows the error, measured in meters or degrees, across all 10 trials for each alignment parameter (rotation and translation). The initializations are represented in blue, indicating large but random initial errors, while the final errors after co-registration are shown in orange, demonstrating successful reduction in alignment error for all six parameters. The overall RMSE values for each co-registration parameter, computed across all scan pairs and initializations, are summarized in the second row of the legend within each plot. We found that the RMSE is less than $0.2^{\circ}$ for rotation and less than $0.045$ meters (m) for translation parameters. These results confirm that the method is flexible and robust, capable of handling vegetation diversity, topographical variations, and random alignment initializations effectively.
@@ 3.2.2 Non-uniform ground terrain
Fig. 6 provides a summary of our experiments evaluating registration performance as a function of percent overlap. Each plot corresponds to an alignment parameter, with blue and orange shapes representing the distributions of initialization and estimate errors relative to ground truth, respectively. The width of each distribution indicates the frequency distribution of error values (vertical axis), while the horizontal axis shows the percent overlap tested across four settings: $28,15,5,1\%$. A total of 1000 co-registration trials per percent overlap were conducted, resulting in 4000 co-registrations in total. For each trial, we randomly initialized the six alignment parameters (simultaneously randomized) from a uniform distribution centered at the true alignment, deviating by at most ±$\pm 45^{\circ}$ in rotation and ±$\pm 15$ m in translation (as shown on the vertical axis). In Fig.6, it is evident that errors for all six parameters remain concentrated around zero, regardless of initialization and the percent overlap tested. Table 1 summarizes the quantitative RMSE values, providing a quantification of this error variability across different alignment parameters and percent overlaps. The consistently small RMSE values across varying levels of overlap underscore the effectiveness and robustness of our approach, even under conditions of limited point-cloud overlap.
@@
Here, we include a performance comparison using a benchmark dataset from [46]. Such dataset was collected with Styrofoam sphere targets manually and strategically placed in the scene. This dataset consists of scans from 17 plots, each with 4 scans per plot arranged in a triangular pattern (one at the center and three at each corner 15 meters apart from the center), along with 9 manually placed targets within the scene.
Comparisons include the manual co-registration approach that uses the CloudCompare software of [44], the automatic and targetless forest registration (TFR) method [25] , the tree-locations based method of [47], and the canopy-based method of [34]. The TFR method from [25] is a trunk-based approach that identifies tree correspondences using the tree diameter at breast height (DBH) biometric and uses these to find correspondences and determine the co-registration parameters between scans. This method is commonly used as a benchmark for forest co-registration comparisons, particularly for TLS-to-TLS co-registration. The tree locations based method extracts the location of each tree from tree trunk information and uses relative positional information to find correspondences between scans. The canopy-based approach of [34] instead segments the tree canopies and extracts features from such segments to find correspondences between scans. In general, these methods were specifically designed for co-registering LiDAR scans from forests, relying in general on tree detection processing, and the use of these features to find correspondences between scans. We conducted a total of 100 trials per available scan pair, initializing each trial uniformly with random rotations and translations within at most ±$\pm 45^{\circ}$ in rotation and ±$\pm 15$ m in translation error relative to ground truth. Table 2 summarizes the RMSE for all 51 pairs of scans across 100 trials each, relative to the ground truth registration parameters already provided in this case by [46]. Throughout all comparative performance results, those highlighted in bold represent the best performance for a given parameter, whereas the second best are highlighted in blue.
Table 2: Registration RMSE relative to benchmark [46] target-based method in TLS-to-TLS
@@
CloudCompare [44] 0.0623 0.0462 0.041 0.031 0.033 0.019
Trunk-based TFR [25] 0.071 0.085 0.092 0.042 0.047 0.022
Tree-locations [47] 0.083 0.044 0.02 0.091 0.085 0.023
Canopy-based [34] 0.322 0.621 0.432 0.153 0.234 0.191
ForestAlign 0.048 0.055 0.032 0.036 0.032 0.007
From Table 2, we see that the TFR and the tree locations method performed similarly, with both slightly better than the canopy based method. This is because both are tree trunk based, and TLS scans are known to have overall good spatial resolution below the canopy. In contrast, the canopy-based method performed decently but not as good as those that are trunk based. Main reason is that TLS is more susceptible to foliage occlusion, and such occlusion varies depending on the scan view-point. Such variation results in varying canopy estimates over scans, which makes finding feature correspondences a bit more challenging for this specific TLS-to-TLS co-registration case. Instead, the proposed approach, which leverages the entire point cloud, including the ground topography for co-registration as opposed to only tree features, performs overall better than all the tree feature based methods. It compares favorably against the manual co-registration method and achieves small error relative to the ground truth parameters provided by the benchmark dataset, and obtained using a reference based method. In our case, however, with the additional benefits of being fully automatic, eliminating the need for manually placed targets in the scene.
Additional experiments were conducted to compare the performance of our co-registration method against standard methods. The goal of adding comparisons with standard fully automated methods was to demonstrate that these methods do not perform well in forest environments. First, we provide further analysis to show the error behavior of the most standard co-registration method (i.e., iterative closest point (ICP)) as initialization deviations increase in forest settings. While ICP is known to achieve outstanding performance in point cloud co-registration for urban scenes, this is not the case in forest point clouds, particularly in the absence of man-made components. For the range of initializations studied in this paper (e.g., within $\pm$10-15 meters in translation $\pm$45o in rotation from ground truth), ICP’s performance in forest scans is less effective. To illustrate this, we conducted experiments using scans from forests in the NM dataset. We co-registered 15 scan pairs randomly selected from the dataset. Each scan was initialized according to a uniform distribution over a range of deviations from ground truth. Co-registration was performed with 50 trials per scan pair and initialization deviation. The results are summarized in Figure 7. The two plots show performance error with respect to ground truth in rotation and translation parameters on the vertical axis, as a function of initialization deviation on the horizontal axis. The plots reveal that ICP performs well within a small initialization deviation from ground truth but deteriorates as deviations increase. This issue is particularly evident for yaw rotation and $x$ and $y$ translations, where performance is more sensitive to larger deviations. ICP struggled with point clouds when translations exceeded 5 meters in translation or yaw rotations exceeded 10o from ground truth. In forest monitoring and surveying using TLS, standard sampling practices often place scans more than these distances apart to enhance spatial coverage and reduce occlusion. As such, standard ICP is not well-suited to the deviations typical of forest sampling protocols. The challenges with ICP in these scenarios are primarily due to the complex structure and spatial randomness of foliage, which can lead to multiple local minima in the least squares formulation of ICP, and which consequently causes ICP to get stuck at the incorrect parameters implied by the multiple local minima. Such an effect becomes more problematic at increasing levels of forest vegetation density, which in turn increase the structural complexity of the scene. Our approach was inspired by such ICP error behavior, and we aimed to address it by first using 3D points from low-complexity structures to obtain a rough alignment, and then incorporating points from higher-complexity structures (e.g., foliage) later, when the estimates are expected to be closer to the ground truth.
(a) Rotation initial deviation
(b) Translation initial deviation
Figure 7: Problem with standard ICP [48] for co-registering point clouds from forests as a function of deviation from ground truth.
@@
Second, we performed a more comprehensive comparison of automatic, targetless methods that do not rely on lines, contours, planes, normals, or edge features, which are typically assumed by methods designed for urban scenarios. The non-forest specific methods we compare include iterative closest point [48], coherent point drift [49], gaussian mixture models [50], support vector registration [51], trees of gaussian mixtures [52], filterReg [53], and Bayesian coherent point drift [54], all available in the Probreg Python package [55]. The forest specific methods we compare are the DBH trunk based TFR method from [25], the tree locations from [47] and the canopy-based method from [34]. In this case, we randomly selected 50 pairs of point clouds from the available forest plots in the NM dataset. Each pair was tested using 100 trials with random initialization from the ground truth co-registration parameters, deviating by ±$\pm 30$ degrees and ±$\pm 15$ meters in a uniform distribution. The RMSE relative to the six degrees of freedom was then computed and summarized in Table 3. The first block of rows corresponds to standard automatic co-registration methods, the second block to methods related to forest features, and the last block summarizes the performance of our approach.
Table 3: RMSE performance comparisons of TLS-to-TLS co-registration methods.